find a formula for the inverse of the function if it exists
f(x)=x/x-1
Thank you tutors

Answer by stanbon(48535)

f(x)=x/x-1
------------
y = x / (x-1)
Interchange x and y to get:
x = y/(y-1)
Solve for y:
y = xy - x
xy -y = x
y(1-x) = x
y = x/(1-x) is the inverse of y = x(x-1)
=======================
Cheers,
Stan H.